hi. my question is given below:
at 25'C henry's constant for CO2 is 1.67 * 10 to the power 8 Pa.Then at the same temp what amount of CO2 is dissolved in 2.25L sodawater at 2.5 atm?
1
Expert's answer
2018-07-09T07:21:21-0400
K = C/P C = K*P = 1.67*10*8 = 133.6 mol/L; nRT = PV n = PV/RT = 2533000*2.25/(8.314*298*1000) = 2.3 mol.